# Five-loop Anomalous Dimensions of Cubic Scalar Theory from Operator Product Expansion

### v2

@@ Abstract
Recently, an Operator Product Expansion method has been proposed for renormalization in quantum field theories. This method extracts UV divergences from two-point propagator-type integrals, and it is believed to be applicable to generic QFTs. In the present work, we apply the OPE method to the renormalization of scalar theory with cubic interactions in six-dimensions. We have successfully computed the 5-loop corrections to the anomalous dimensions of the $\phi^{Q}$ operator, as well as the large $N$ expansion of scaling dimensions at the Wilson-Fisher fixed point to $1/N^{5}$ order. This computation sets a new loop-order record for the anomalous dimension of the $\phi^{Q}$ operator in cubic scalar theory, and further demonstrates the efficiency and capacity of the OPE method in renormalization.
@@ 1 Introduction
During the development of the renormalization group method, scalar field theories have served as an ideal playground for testing multi-loop computation techniques and providing insights into more complicated field theories such as non-Abelian gauge theories. Yet the renormalization of scalar field theories is itself of great importance, as explored in Wilson’s theory of critical phenomena Wilson:1971dc , that in certain statistical systems, the critical exponents near second-order phase transitions are related to the anomalous dimensions of the field or composite operators constructed from fields PELISSETTO2002549 ; Moshe:2003xn . Two notable scalar field theories are the six-dimensional cubic scalar field theory and the four-dimensional quartic scalar theory, both of which are renormalizable theories. These two theories belong to the same $O(N)$ universality class, and it has been conjectured that the $O(N)$ quartic scalar theory at its UV fixed point is equivalent to the $O(N)$ cubic scalar theory at its IR fixed point Fei:2014yja ; Fei:2014xta , namely they approach to the same critical theory in the fixed points. Verifying this conjecture with high-loop precision for the scaling dimensions of fields or composite operators represents an intriguing problem, but it demands perturbative computation of renormalization functions to higher-loop orders. Because of the technical advancements, recently there has been a growing interest in pushing renormalization computations to the next loop order.
@@
It has been a long history for renormalization computations of quartic and cubic scalar theories, dating back to the early era studies of critical exponents up to $\epsilon^{2}$ orders for generalized Ising and Baxter models Wilson:1971dc , as well as the generalized Heisenberg model Wilson:1971vs . The 3-loop computations of the beta functions, field renormalization and mass renormalization of the quartic scalar theory were completed just one year after the 1-loop and 2-loop results BREZIN1973227 ; Brezin:1973igb , with 4-loop results following five years later Kazakov:1979ik . The full 5-loop renormalization of the quartic scalar theory required over a decade of collaborative efforts from numerous researchers Chetyrkin:1981jq ; Gorishnii:1983gp ; Kazakov:1983dyk ; Gorishnii:1983jr ; Kleinert:1991rg ; Adzhemyan:2013jra . Throughout this development, techniques such as the Gegenbauer polynomial $x$-space, IR-rearrangement Vladimirov:1979zm ; Chetyrkin:1980pr ; Caswell:1981ek , $R$-operation ($R^{\ast}$-operation) Chetyrkin:1982nn ; CHETYRKIN1984419 ; Larin:2002sc , as well as the modern Feynman diagram generating and Feynman integral reduction algorithms were refined and played the major roles. Early attempts were also made to push renormalization computations to higher-loop order, for instance the primitive divergences of the 7-loop beta function Broadhurst:1995km , or extend them to composite operators, for instance the 4-loop anomalous dimension of operators with derivatives Derkachov:1997pf . However, the complete 6-loop computation of the quartic scalar theory became feasible only approximately 15 years after the 5-loop result, firstly with the 6-loop field renormalization result Batkovich:2016jus , followed by the beta function and mass renormalization for the $O(1)$ model Kompaniets:2016hct , and subsequently for the $O(N)$ model Kompaniets:2017yct . A major breakthrough came with the development of the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which enabled the successful computation of the 7-loop beta function, mass renormalization, and field renormalization Schnetz:2016fhy , and even the 8-loop field renormalization function Schnetz:2025opm . The history of renormalization computations for the cubic scalar theory has closely paralleled that of the quartic scalar theory, which is not surprising since both have benefited from the same technical advancements. However, due to greater complexity, progress in the cubic scalar theory has consistently lagged by 1 or 2 loops compared to its quartic counterpart at similar stages. Early 1-loop and 2-loop analysis of the cubic scalar theory revealed its asymptotic freedom Macfarlane:1974vp ; Ma:1975vn . Over the subsequent 40 years, 3-loop deAlcantaraBonfim:1980pe ; deAlcantaraBonfim:1981sy ; Fei:2014xta and 4-loop Gracey:2015tta renormalization computations were gradually completed, with the 5-loop result finished only very recently Kompaniets:2021hwg ; Borinsky:2021jdb . The frontier record once again came from the graphical function method, by which the 6-loop beta function, field renormalization and mass renormalization functions of the cubic scalar theory have been successfully acquired this very year Schnetz:2025wtu .
@@
Besides the computations of the beta functions and anomalous dimensions of fields and mass, the renormalization of composite operators constitutes an equal important yet more challenging problem. Phenomenologically, this is relevant to specific critical exponents in phase transition problems. One class of composite operators of interest is the fixed-charge symmetric traceless operator $\phi^{Q}$. In statistical physics, the anomalous dimensions of $\phi^{Q}$ for certain values of $Q$ are relevant to phase transitions in diverse systems PELISSETTO2002549 , such as liquid crystals, graphite-intercalation compounds, polymers, etc. DePrato:2003yd ; Calabrese:2004ca . Interest in computing the scaling dimensions of the $\phi^{Q}$ operator in cubic and quartic scalar theories grows from 2019, when the semi-classical method is introduced to compute the scaling dimension of $\phi^{Q}$ for generic $Q$ at the large $Q$ limit. This leads to all-loop results at leading and sub-leading orders in the $Q$ expansion for the quartic scalar theory Arias-Tamargo:2019xld ; Badel:2019oxl ; Antipin:2020abu ; Giombi:2020enj and the cubic scalar theory Arias-Tamargo:2020fow ; Antipin:2021jiw . Progress in the semi-classical approach stimulates the growing attention to the perturbative method approach towards renormalization computation of the $\phi^{Q}$ operator, aiming to validate the semi-classical results and also to obtain results for all orders of $Q$. For the quartic scalar theory, the scaling dimension of the $\phi^{Q}$ operator has been computed perturbatively at 4-loop Jack:2021ypd and 5-loop order Jin:2022nqq . While the 6-loop result is also available Bednyakov:2022guj , in which the data of $Q=1,\ldots,5$ was computed via perturbative Feynman diagram method and the data of $Q=6,7$ was imported from the semi-classical method. Concurrently, for the cubic scalar theory, perturbative computations have been carried out at 3-loop Jack:2021ziq and 4-loop order Huang:2024hsn .
High-loop order renormalization computations would never be possible without the continuous evolution of both traditional and modern techniques. One of the important techniques to mention is the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which takes the advantage of computing scalar integrals in coordinate space. It pushes renormalization computations to 8-loops for the four-dimensional quartic scalar theory and 6-loops for the six-dimensional cubic scalar theory, and for the first time, with some tricks it also pushes the renormalization computations to 7-loops in the Gross-Neveu model involving spinning particles Huang:2025ree . The graphical function method is designed to evaluate scalar integrals with no more than four external points, which is certainly not the case for integrals of correlation functions involving $\phi^{Q}$ operators. To address this, we have proposed an Operator Product Expansion (OPE) method for renormalization computations by analyzing the UV divergence structures of correlation functions in the large momentum expansion limit Huang:2024hsn . This OPE method has proven highly useful and efficient for determining anomalous dimensions Huang:2024hsn , and this fact has also been shown in the condensate calculations Marino:2024uco ; Liu:2025bqq . In the OPE framework, the momenta of the two operators are taken to the infinite limit. This can be considered pictorially as two hard fields propagating in a soft field background Huang:2016bmv ; Huang:2022qnx , reducing the integral evaluation to a single-scale problem characterized by the large momentum parameter. By identifying the contributing terms in the OPE, the essential UV divergences required for renormalization computations are found to be determined by two-point hard sub-graphs. Concurrently, the Wilson coefficients in the OPE are proportional to a combination of $Z$-factors and the contributions of these two-point hard sub-graphs. By requiring the UV finiteness of the Wilson coefficients, the $Z$-factors can be determined once the UV divergences of the hard graphs are known. This approach thus reduces renormalization computations to the evaluation of two-point propagator-type integrals. Note that the generality of OPE across any quantum field theories enables diverse applications of this OPE method, to theories with scalars or particles with spins, and also to the situations involving operator mixing and operators with derivatives, with Lorentz indices, etc.
@@
In this work, we apply the OPE method developed in Huang:2024hsn to compute the 5-loop scaling dimension of the $\phi^{Q}$ operators in the six-dimensional $O(N)$ cubic scalar theory, thereby extending beyond the 4-loop result reported in that paper. This paper is structured as follows. In §2, we provide a concise review of the cubic scalar theory, its renormalization functions, and the OPE method for renormalization computation. Treatments for organizing and simplifying Feynman graphs of the correlation functions involving the $\phi^{Q}$ operator are also presented. The §3 is devoted to a detailed demonstration of the OPE method using 2-loop examples, showing how the field renormalization function, beta function and the anomalous dimensions of the $\phi^{Q}$ operator can be computed efficiently. In §4, general discussion for the computation of 5-loop anomalous dimension of $\phi^{Q}$ operator has been presented, and various 5-loop results can be found in this section and the Appendix. In §5, conclusion and discussion are presented.
@@ 2 Review of the theory and the method
In this section we review the six-dimensional $O(N)$ cubic scalar theory, with emphasis on its renormalization computations and the techniques for organizing and simplifying Feynman diagrams. To make this paper self-contained, we also provide a brief review of the OPE method for renormalization computations Huang:2024hsn , especially for its application in the context of cubic scalar theory.
@@ 2.1 The cubic scalar theory and its renormalization
$\mathcal{L}=\frac{1}{2}(\partial\phi_{i})^{2}+\frac{1}{2}(\partial\sigma)^{2}-\frac{g}{2!}\sigma\phi_{i}^{2}-\frac{h}{3!}\sigma^{3}~,$ (1)
where $\sigma$ and $\phi_{i},i=1,\ldots,N$ denote two types of real scalar fields, with the $\phi_{i}$ fields transforming as an $O(N)$ vector. This theory is perturbatively renormalizable in six-dimensional space-time. The primary focus of this paper is the renormalization analysis of the symmetric traceless $\phi^{Q}$ operator, defined as,
@@
where $T_{i_{1}\cdots i_{Q}}$ is a symmetric traceless tensor and $Q$ is a fixed charge number. We will work in the $\overline{\mbox{MS}}$ regularization scheme tHooft:1973mfk in $d=(6-\epsilon)$-dimension. The classical scaling dimension of $\phi^{Q}$ is Q(d/21)Q(d/2-1), while the total scaling dimension is
$\Delta_{Q}=Q\left(\frac{d}{2}-1\right)+\gamma_{Q}+Q\gamma_{\phi}~.$ (3)
The difference between the total scaling dimension and the classical scaling dimension is defined as the anomalous dimension of the $\phi^{Q}$ operator. The gamma function $\gamma_{Q}$ of $\phi^{Q}$ operator can be computed perturbatively by Feynman diagrams of inserting the $\phi^{Q}$ operator in the 1PI Green function with $Q$ external legs. However, explicit evaluation of these correlation functions is complicated by the cumbersome $O(N)$ tensor structures inherent to the operator, which cause the number of terms in Feynman rules to grow exponentially with increasing $Q$. To address this problem, it is convenient to introduce an auxiliary $O(N)$ zero-norm complex vector $q$ satisfying $^{2}=0, qq$$\cdot\bar{q}=1$, and define the tensor $T$ as Antipin:2020abu ; Jin:2022nqq
@@
This allows us to redefine the $\phi^{Q}$ operator in terms of a $\varphi^{Q}$ operator,
@@
where the complex field $\varphi$ is a redefinition of the scalar field $\varphi:=q\cdot\phi$, with its complex conjugate field $\bar{\varphi}:=\bar{q}\cdot\phi$. This definition effectively strips $O(N)$ indices from the correlation functions, while the anomalous dimension of $\phi^{Q}$ operator is equivalent to that of $\varphi^{Q}$ operator. The latter can now be computed by $\varphi^{Q}$-$Q\varphi$ correlation functions, which are free of $O(N)$ index structures. In the $\varphi$-field representation, the $\sigma\phi^{2}$ vertices along the line of the operator and the external legs are reformulated as $\sigma\varphi\bar{\varphi}$ vertices, while other $\sigma\phi^{2}$ vertices appearing in loop structures of $\varphi$ fields contribute factors of $N$ after summing over color indices. Due to charge conservation, scalar charge flows along $\varphi$ fields in the corresponding Feynman diagrams.
For purpose of renormalization computation, following standard procedures, we introduce renormalized fields as,
@@
and renormalized coupling constants as,
$g_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{g}g~~~,~~~h_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{h}h~~~\mbox{with}~~~\widetilde{Z}_{g}=Z_{g}Z^{-\frac{1}{2}}_{\sigma}Z^{-1}_{\varphi}~~~,~~~\widetilde{Z}_{h}=Z_{h}Z^{-\frac{3}{2}}_{\sigma}~,$ (7)
where $g_{0},h_{0}$ denote bare coupling constants and $\varphi_{0},\sigma_{0}$ denote bare fields. We also define the renormalized operator $\varphi^{Q}_{R}=Z_{\varphi^{Q}}\varphi_{0}^{Q}$. Since $\varphi_{0}^{Q}=Z_{\varphi}^{\frac{Q}{2}}\varphi^{Q}$, we get
@@
Then $\widetilde{Z}_{\varphi^{Q}}$ can be assigned as the Feynman rule of the $\varphi_{R}^{Q}$-$Q\varphi$ vertex, and the anomalous dimension of the $\varphi^{Q}$ operator can be determined by the $Z_{\varphi^{Q}}$-factor.
Provided that the $Z_{\varphi}$, $Z_{\sigma}$, $Z_{g}$ and $Z_{h}$ have been determined, we can compute beta functions via the standard field theory method. The beta functions $\beta(g)=\frac{\partial g}{\partial\ln\mu}$ and $\beta(h)=\frac{\partial h}{\partial\ln\mu}$ for the renormalized couplings can be derived from $\widetilde{Z}_{h},\widetilde{Z}_{g}$ by solving the following matrix,
@@
Then the field anomalous dimensions of $\varphi$ and $\sigma$ can be computed by definition as,
@@
Furthermore, once $Z_{\varphi^{Q}}$ has been determined, the anomalous dimension of the $\varphi^{Q}$ operator can be computed by
$\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}=-\frac{\partial\ln Z_{\varphi^{Q}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}Z_{\varphi}^{-\frac{Q}{2}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}+Q\left(\frac{1}{2}\frac{\partial\ln Z_{\varphi}}{\partial\ln\mu}\right)~,$ (11)
where the first term computes $\gamma_{Q}:=\gamma_{\varphi^{Q}}$ and the second term computes $Q\gamma_{\varphi}$.
With the knowledge of beta functions and the anomalous dimensions, in order to study the large $N$ expansion of the anomalous dimensions around the Wilson-Fisher fixed point, one need to solve the fixed points from beta functions,
@@
After substituting the solutions $g^{\ast},h^{\ast}$ into the anomalous dimensions and expanding the results in the series of $1/N$, one produces the large $N$ expansion of the cubic scalar theory at the IR fixed point.
The renormalization computation problems require the determination of the $Z$-factors. In general, a $Z$-factor of up to $L$-loop contributions takes the schematic form as

### v4

@@ Abstract
In this work, we compute the anomalous dimensions of the $\phi^{Q}$ operator in six-dimensional cubic scalar theory. The renormalization analysis is carried out within the framework of the Operator Product Expansion method, while the ultraviolet divergences of Feynman integrals are evaluated using the graphical function method. Inspired by the intrinsic connection between Wilson coefficients and anomalous dimensions, an algorithm was proposed recently, which provides a practical and systematic framework for calculating the anomalous dimensions of masses, fields, and composite operators, with broad potential applicability to generic quantum field theories. Notably, the HyperlogProcedures package, developed based on the graphical function method, enables the computation of two-point propagator-type integrals, derived herein for capturing ultraviolet divergences, to very high loop orders. With these advanced techniques, we have successfully computed the anomalous dimensions of the $\phi^{Q}$ operator up to five loops. Furthermore, we present a large $N$ expansion of the scaling dimensions at the Wilson-Fisher fixed point, extended to the $1/N^{5}$ order. This computation sets a new loop-order record for the anomalous dimension of the $\phi^{Q}$ operator in cubic scalar theory, while further validating the efficiency and versatility of the proposed algorithm in renormalization analyses.
@@ 1 Introduction
In the development of the renormalization group (RG) method, scalar field theories have emerged as an ideal testing ground for validating multi-loop computation techniques while offering valuable insights into more intricate field theories, such as non-Abelian gauge theories. Nevertheless, the renormalization of scalar field theories is inherently significant in its own right. As elucidated in Wilson’s theory of critical phenomena Wilson:1971dc , critical exponents near second-order phase transitions in certain statistical systems are intimately linked to the anomalous dimensions of fields or composite operators PELISSETTO2002549 ; Moshe:2003xn . Two prominent examples of renormalizable scalar field theories are the six-dimensional cubic scalar field theory and the four-dimensional quartic scalar field theory. Both theories belong to the same $O(N)$ universality class, and a key conjecture posits that the $O(N)$ quartic scalar theory at its ultraviolet (UV) fixed point is equivalent to the $O(N)$ cubic scalar theory at its infrared (IR) fixed point Fei:2014yja ; Fei:2014xta . Namely, they approach to the same critical theory at their respective fixed points. Verifying this conjecture with high-loop precision for the scaling dimensions of fields or composite operators represents an intriguing problem, yet it requires perturbative computations of renormalization functions to higher-loop orders. Thanks to recent technical advancements, there has been a growing surge of interest in pushing renormalization computations to the next loop order.
@@
The renormalization computations of quartic and cubic scalar theories possess a long-standing history, tracing back to early-era studies of critical exponents up to $\epsilon^{2}$ orders for generalized Ising and Baxter models Wilson:1971dc , as well as the generalized Heisenberg model Wilson:1971vs . The 3-loop computations of the beta functions, field renormalization and mass renormalization for the quartic scalar theory were completed just one year later BREZIN1973227 ; Brezin:1973igb , with 4-loop results following five years thereafter Kazakov:1979ik . The full 5-loop renormalization of the quartic scalar theory, however, demanded over a decade of collaborative efforts from numerous researchers Chetyrkin:1981jq ; Gorishnii:1983gp ; Kazakov:1983dyk ; Gorishnii:1983jr ; Kleinert:1991rg ; Adzhemyan:2013jra . Throughout this development, techniques such as the IR-rearrangement Vladimirov:1979zm ; Chetyrkin:1980pr ; Caswell:1981ek , the $R$-operation ($R^{\ast}$-operation) Chetyrkin:1982nn ; CHETYRKIN1984419 ; Larin:2002sc , and modern Feynman diagram generating and Feynman integral reduction algorithms were refined and played major roles. Early attempts were also made to push renormalization computations to higher-loop orders, for instance the primitive divergences of the 7-loop beta function Broadhurst:1995km . Or extend them to composite operators, for instance the 4-loop anomalous dimension of operators with derivatives Derkachov:1997pf . Nevertheless, the complete 6-loop computation of the quartic scalar theory only became feasible approximately 15 years after the 5-loop result. Firstly with the 6-loop field renormalization result Batkovich:2016jus , followed by the beta function and mass renormalization for the $O(1)$ model Kompaniets:2016hct , and subsequently for the $O(N)$ model Kompaniets:2017yct . A major breakthrough emerged with the development of the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which enabled the successful computation of the 7-loop beta function, mass renormalization, and field renormalization Schnetz:2016fhy , as well as the 8-loop field renormalization function Schnetz:2025opm . The history of renormalization computations for the cubic scalar theory has closely paralleled that of the quartic scalar theory, which is not surprising since both have benefited from the same technical advancements. However, due to its greater complexity, progress in the cubic scalar theory has consistently lagged by 1 or 2 loops compared to its quartic counterpart at similar stages. Early 1-loop and 2-loop analysis of the cubic scalar theory revealed its asymptotic freedom Macfarlane:1974vp ; Ma:1975vn . Over the subsequent 40 years, 3-loop deAlcantaraBonfim:1980pe ; deAlcantaraBonfim:1981sy ; Fei:2014xta and 4-loop Gracey:2015tta renormalization computations were gradually completed, with the 5-loop result finished only very recently Kompaniets:2021hwg ; Borinsky:2021jdb . Once again, the frontier record stems from the graphical function method, through which the 6-loop beta function, field renormalization and mass renormalization functions of the cubic scalar theory have been successfully acquired this year Schnetz:2025wtu .
@@
Besides the beta functions and the anomalous dimensions of fields and mass, the renormalization of composite operators constitutes an equally important yet more challenging problem. Phenomenologically, this is relevant to specific critical exponents in phase transition phenomena. One class of composite operators of interest is the fixed-charge symmetric traceless operator $\phi^{Q}$. In statistical physics, the anomalous dimensions of $\phi^{Q}$ for certain values of $Q$ are relevant to phase transitions in diverse systems PELISSETTO2002549 , such as liquid crystals, graphite-intercalation compounds and polymers DePrato:2003yd ; Calabrese:2004ca . Interest in computing the scaling dimensions of the $\phi^{Q}$ operator in cubic and quartic scalar theories has grown since 2019, when the semi-classical method is introduced to compute the scaling dimension of $\phi^{Q}$ at the large $Q$ limit. This advancement leads to all-loop results at leading and sub-leading orders in the $Q$ expansion for the quartic scalar theory Arias-Tamargo:2019xld ; Badel:2019oxl ; Antipin:2020abu ; Giombi:2020enj and the cubic scalar theory Arias-Tamargo:2020fow ; Antipin:2021jiw . Progress in the semi-classical approach has stimulated increasing attention to the perturbative method approach towards renormalization computation of the $\phi^{Q}$ operator, aiming to validate the semi-classical results and also to obtain results for complete $Q$ dependence. For the quartic scalar theory, the scaling dimension of the $\phi^{Q}$ operator has been computed perturbatively at 4-loop Jack:2021ypd and 5-loop orders Jin:2022nqq . While the 6-loop result is also available Bednyakov:2022guj , where the data of $Q=1,\ldots,5$ was computed via perturbative Feynman diagram method and the data of $Q=6,7$ was adopted from the semi-classical method. Concurrently, for the cubic scalar theory, perturbative computations have been carried out at 3-loop Jack:2021ziq and 4-loop orders Huang:2024hsn .
High-loop order renormalization computations would never be feasible without the continuous evolution of both traditional and modern techniques. One notable approach is the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which takes the advantage of computing scalar integrals in coordinate space. To date, it has advanced renormalization computations to 8-loop order in the four-dimensional quartic scalar theory and 6-loop order in the six-dimensional cubic scalar theory. Furthermore, with some tricks, it has for the first time pushed renormalization computations to 7-loop order in the Gross-Neveu model Huang:2025ree . However, the graphical function method is inherently limited to scalar integrals with no more than four external points. When investigating the anomalous dimensions of higher-length composite operators, for instance the $\phi^{Q}$ operator, the number of external legs can be very large, rendering the graphical function method inapplicable directly. To address this challenge, a novel renormalization algorithm based on the Operator Product Expansion (OPE) has been proposed Huang:2024hsn . As established in the literature, the renormalization group equation (RGE) flow of Wilson coefficients is governed by the anomalous dimensions of operators, e.g., as described in Collins_1984 . This principle enables the extraction of anomalous dimensions from Wilson coefficients and has been occasionally employed to compute the anomalous dimensions of specific operators. For instance, the OPE of two scalar fields was used to determine the anomalous dimension of the $\phi^{2}$ operator Collins_1984 , and similarly, the OPE of two stress-energy tensors was applied to compute the anomalous dimension of the Konishi operator in $\mathcal{N}=4$ super-Yang-Mills theory Eden:2012fe . In practice, however, standard renormalization computations typically rely on traditional techniques such as the $R^{\ast}$-operation method and the massive vacuum bubble method. By considering the OPE of a composite operator with a fundamental field, the proposed algorithm Huang:2024hsn extends the applicability of the OPE framework to general composite operators in a systematic manner, making it suitable for implementation in modern computational tools. In this algorithm, the UV finiteness of the corresponding OPE coefficients links the anomalous dimensions of higher-dimensional composite operators to those of lower-dimensional operators, thereby establishing recursive relations for anomalous dimensions. Utilizing these recursion relations, the anomalous dimensions of complicated operators can be non-trivially derived from those of simple operators. Additionally, the UV divergences required for renormalization computation are treated globally without the need to subtract sub-divergences, and the latter is often computationally cumbersome in other renormalization techniques such as the $R^{\ast}$-operation method. The algorithm also provides a systematic approach to deform integrals of correlation functions involving fields and/or operators into two-point propagator-type integrals, which is particularly advantageous when dealing with systems with numerous external legs. We have explicitly validated the OPE based algorithm in the cubic scalar theory and Gross-Neveu-Yukawa theory Huang:2024hsn , confirming its highly versatile and efficient. The utility of determining anomalous dimensions via OPE has also been tested in condensate calculations Marino:2024uco ; Liu:2025bqq . Notably, the generality of OPE across quantum field theories allows the proposed algorithm to be applied to diverse scenarios, ranging from theories involving scalars or particles with spins to situations that include operator mixing, as well as operators carrying derivatives and Lorentz indices.
@@
In this study, we apply the OPE based algorithm proposed in Huang:2024hsn , combined with the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , to calculate the 5-loop scaling dimension of the $\phi^{Q}$ operators in the six-dimensional $O(N)$ cubic scalar theory, thereby extending beyond the 4-loop result. This paper is structured as follows. In §2, we provide a concise review of the cubic scalar theory, its renormalization functions, and the aforementioned renormalization algorithm. Additionally, we present approaches for organizing and simplifying Feynman graphs of correlation functions involving the $\phi^{Q}$ operator. §3 is dedicated to a detailed demonstration of the OPE based algorithm via 2-loop examples, illustrating how the field renormalization function, beta function, and anomalous dimensions of the $\phi^{Q}$ operator can be computed efficiently. In §4, we present a general discussion on the computation of the 5-loop anomalous dimension of the $\phi^{Q}$ operator, with various 5-loop results provided in this section and the Appendix. In §5, conclusion and discussion are presented.
@@ 2 Review of the theory and the method
In this section, we review the six-dimensional $O(N)$ cubic scalar theory, with a focus on its renormalization computations and the techniques for organizing and simplifying Feynman diagrams. To make this paper self-contained, we also present a brief overview of the OPE based renormalization algorithm Huang:2024hsn , particular in the context of its application to the cubic scalar theory.
@@ 2.1 The cubic scalar theory and its renormalization
$\mathcal{L}=\frac{1}{2}(\partial\phi_{i})^{2}+\frac{1}{2}(\partial\sigma)^{2}+\frac{g}{2!}\sigma\phi_{i}^{2}+\frac{h}{3!}\sigma^{3}~,$ (1)
where $\sigma$ and $\phi_{i},i=1,\ldots,N$ denote two types of real scalar fields, with the $\phi_{i}$ transforming as an $O(N)$ vector. This theory is perturbatively renormalizable in six-dimensional spacetime. The primary focus of this paper is the renormalization analysis of the symmetric traceless $\phi^{Q}$ operator, defined as,
@@
where $T_{i_{1}\cdots i_{Q}}$ is a symmetric traceless tensor and $Q$ is a fixed charge number. We work in the $\overline{\mbox{MS}}$ regularization scheme tHooft:1973mfk in $d=(6-\epsilon)$-dimension. The classical scaling dimension of $\phi^{Q}$ is Q(d/21)Q(d/2-1), while the total scaling dimension reads
$\Delta_{Q}=Q\left(\frac{d}{2}-1\right)+\gamma_{\varphi^{Q}}~.$ (3)
The difference between the total and classical scaling dimensions defines the anomalous dimension $\gamma_{\phi^{Q}}$ of the $\phi^{Q}$ operator. It can be computed perturbatively via Feynman diagrams involving the insertion of the $\phi^{Q}$ operator into 1PI Green functions with $Q$ external legs. However, explicit evaluation of these correlation functions is complicated by the cumbersome $O(N)$ tensor structures inherent to the operator, which cause the number of terms in the Feynman rules to grow exponentially with increasing $Q$. To address this challenge, it is convenient to introduce an auxiliary $O(N)$ zero-norm complex vector $q$ satisfying $^{2}=0, qq$$\cdot\bar{q}=1$, and define the tensor $T$ as Antipin:2020abu ; Jin:2022nqq
@@
This allows us to reexpress the $\phi^{Q}$ operator in terms of a $\varphi^{Q}$ operator,
@@
where the complex field $\varphi$ is defined as $\varphi:=q\cdot\phi$, with its complex conjugate $\bar{\varphi}:=\bar{q}\cdot\phi$. This redefinition effectively eliminates $O(N)$ indices from the correlation functions, and the anomalous dimension of $\phi^{Q}$ operator is equivalent to that of $\varphi^{Q}$ operator. The latter can now be computed by $\varphi^{Q}$-$Q\varphi$ correlation functions, which are free of $O(N)$ index structures. In the $\varphi$-field representation, the $\sigma\phi^{2}$ vertices along the operator and external legs are reformulated as $\sigma\varphi\bar{\varphi}$ vertices, while other $\sigma\phi^{2}$ vertices appearing in loop structures of $\varphi$ fields contribute factors of $N$ after summing over $O(N)$ indices. Owing to charge conservation, scalar charge flows along the $\varphi$ fields in the corresponding Feynman diagrams.
For the purpose of renormalization computation, we follow standard textbook procedures and introduce renormalized fields as
@@
along with renormalized coupling constants as
$g_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{g}g~~~,~~~h_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{h}h~~~\mbox{with}~~~\widetilde{Z}_{g}=Z_{g}Z^{-\frac{1}{2}}_{\sigma}Z^{-1}_{\varphi}~~~,~~~\widetilde{Z}_{h}=Z_{h}Z^{-\frac{3}{2}}_{\sigma}~.$ (7)
Here, $g_{0},h_{0}$ denote bare coupling constants and $\varphi_{0},\sigma_{0}$ denote bare fields. We also define the renormalized operator $\varphi^{Q}_{R}=Z_{\varphi^{Q}}\varphi_{0}^{Q}$, where the anomalous dimension $\gamma_{\varphi^{Q}}$ is derived from $Z_{\varphi^{Q}}$. Since $\varphi_{0}^{Q}=Z_{\varphi}^{\frac{Q}{2}}\varphi^{Q}$, we obtain
@@
The factor $\widetilde{Z}_{\varphi^{Q}}$ can then be assigned as the Feynman rule for the $\varphi_{R}^{Q}$-$Q\varphi$ vertex.
Once $Z_{\varphi}$, $Z_{\sigma}$, $Z_{g}$ and $Z_{h}$ are determined, beta functions can be computed using standard field theory methods. These beta functions follow from $\frac{\partial g}{\partial\ln\mu}$ and $\frac{\partial h}{\partial\ln\mu}$, which are derived by solving matrix equation,
@@
The field anomalous dimensions of $\varphi$ and $\sigma$ are then computed by definition as
@@
Furthermore, once $Z_{\varphi^{Q}}$ is determined, the anomalous dimension of the $\varphi^{Q}$ operator is given by
$\gamma_{\varphi^{Q}}:=-\frac{\partial\ln Z_{\varphi^{Q}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}Z_{\varphi}^{-\frac{Q}{2}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}+Q\left(\frac{1}{2}\frac{\partial\ln Z_{\varphi}}{\partial\ln\mu}\right)~.$ (11)
We denote the first term as $\gamma_{Q}$, and the second term evaluates to $Q\gamma_{\varphi}$, such that $\gamma_{\varphi^{Q}}=\gamma_{Q}+Q\gamma_{\varphi}$.
To study the large $N$ expansion of the anomalous dimensions around the Wilson-Fisher fixed point, we first solve for the fixed points of the beta functions,
@@
Substituting the solutions $g^{\ast},h^{\ast}$ into the anomalous dimensions and expanding the results in a series of $1/N$ yields the large $N$ expansion of the cubic scalar theory at the IR fixed point.
Renormalization computations require determining the $Z$-factors. In general, a $Z$-factor including up to $L$-loop contributions takes the schematic form